OpenStudy (anonymous):

Differential Equations HELP! dx/dt-x^3=x I separated them dx-x^3/x=dt not sure if this is right

OpenStudy (zzr0ck3r):

\[\frac{dx}{dt}-x^3=x?\]

OpenStudy (zzr0ck3r):

or \[\frac{dx}{dt}*(-x^3)=x\]

OpenStudy (zzr0ck3r):

@rperez36 ?

OpenStudy (zzr0ck3r):

\[\frac{dx}{dt}=x+x^3=x(1+x^2)\] \[\frac{1}{x(1+x^2)}dx=dt\]

OpenStudy (zzr0ck3r):

can you finish it now?

OpenStudy (anonymous):

yes I will give it a shot. thanks for helping me with my obvious mistake of trying to divide out the -x^3

OpenStudy (anonymous):

just been a long day I should probably get to sleep

OpenStudy (zzr0ck3r):

all good:) np